what is the answer when you have to solve for W when the problem is 2w-y=7w-2
Found 2 solutions by solver91311, tommyt3rd:
Answer by solver91311(24713)



You cannot solve for when the problem is

There is no in because and are NOT the same thing.

You can, however solve for :

Add to both sides

Add to both sides

Collect like terms.

Multiply by the reciprocal of the resulting coefficient on .

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by tommyt3rd(5050)

2w-y=7w-2
7w-2w=2-y
5w=2-y
w=(2-y)/5